A store sells 90 ski jackets for $200 dollars each. A survey indicates that for each $10 decrease in price, five more jackets will be sold.
a) Find the number of jackets sold and the selling price, when the total revenue is $17 600
b) What is the lowest price the will give revenues of at least 15 600? How many jackets would be sold at this price?

Answer by Aldorozos(172)

(90+5x)(200-10x) = 17600
Therefore
-50x^2 +100x +1800 = 17600
We divide both sides by -50
x^2 -2x -8 = 0 This is quadratic equation x = 4. This means that we have dropped the price four times. Which means 4 *10 = 40 Therefore the price = 200-40 = 160 and we have sold 4*5 + 90 jackets which is 110
To see if we have done it correctly we multiply 110* 160 and we get 17600
We solve the second part of the problem the same way and we get x=8
This means we have dropped the price 8 times which is 8*10 = 80. This means the price has to be 200 - 80 = 120
and the number of jackets should be 90+ 8*5 = 130